OpenStudy (anonymous):

Sulfur 4 naturally occurring stable isotopes. Use the relative abundance of each to calculate the atomic mass of sulfur. 95.02% of S with mass number 32, 0.75% of S with mass number 33, 4.21% of S with mass number 34 and 0.02% of sulfur with mass number 36.

OpenStudy (gebooors):

Relative mass of natural Sulphur is (95.02 / 100) * 32 + (0.75/100) * 33 + (4.21/100) * 34 +(0.02 / 100) * 36 I think result is close 32.07, but e.g. for isotope 32 you find from tables more accurate value 31.972 070

OpenStudy (jebonna):

All you have to do for these types of questions, is times the relative abundance by the mass number, add them all together and divide by 100. \[\frac{ (abundance \times mass number) + (abundance \times mass number) }{ 100 }\] Do this for however many isotopes you have in your question. For your asked question you will have to do (abundance x mass number) 4 times as there are 4 different isotopes. I hope this helps!